Iron (III) chloride reacts with magnesium oxide. If 13.75 g of iron (III) chloride reacts with magnesium oxide to form 5.25 g of the magnesium product, what is the percent yield?


1
Expert's answer
2021-03-12T06:00:52-0500

"2FeCl_3+3MgO=Fe_2O_3+3MgCl_2"

m(MgCl2) = M(MgCl2)×n(MgCl2)

n(MgCl2) = 3/2 n(FeCl3) = 3/2 m(FeCl3)/M(FeCl3)

m(MgCl2) = 3/2 M(MgCl2) m(FeCl3)/M(FeCl3) = 12.11 g

Yield = 5.25/12.11 × 100% = 43.35%